OpenStudy (sleepyjess):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 10.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

centered at the origin?

OpenStudy (sleepyjess):

I think so

OpenStudy (jdoe0001):

major axis of 18, is vertical, thus the vertical is the longest axis meaning the "a" component goes under the "y" thus \(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ a}}^2}=1 \quad center\ ({\color{brown}{ 0}},{\color{blue}{ 0}})\quad \cfrac{(x-{\color{brown}{ 0}})^2}{{\color{purple}{ 10}}^2}+\cfrac{(y-{\color{blue}{ 0}})^2}{{\color{purple}{ 18}}^2}=1 \\ \quad \\ \cfrac{x^2}{100}+\cfrac{y^2}{324}=1\)

OpenStudy (jdoe0001):

hmmm actually...,

OpenStudy (jdoe0001):

if the vertical axis is ... lemme draw it up

OpenStudy (jdoe0001):

 thus that means that the "a" component is 9 and "b" is 5

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ 0}},{\color{blue}{ 0}})\implies \cfrac{(x-{\color{brown}{ 0}})^2}{{\color{purple}{ 5}}^2}+\cfrac{(y-{\color{blue}{ 0}})^2}{{\color{purple}{ 9}}^2}=1 \\ \quad \\ \cfrac{x^2}{25}+\cfrac{y^2}{81}=1\)